OpenStudy (anonymous):

whats the surface area of a cylinder with the hieght of 10 cm and a diameter of 10 cm

OpenStudy (lgbasallote):

here's a hint: \[\huge SA = 2\pi (\frac d2)^2 + \pi d h\] where: d = diameter h = height does that help?